OpenStudy (anonymous):

Write the equation of a line that goes through the points (3, -4) and (3, -2)

OpenStudy (anonymous):

First lets find the slope through the points (3, -4) and (3, -2)

OpenStudy (anonymous):

Use the slope formula: \[m = \frac{ y _{2} - y _{1} }{ x _{2} - x _{1} }\]

OpenStudy (anonymous):

@Weech31912 ?

OpenStudy (anonymous):

\[m = \frac{ (-2)- (-4) }{ 3 - 3 }\]

OpenStudy (anonymous):

\[m = \frac{ 2 }{ 0 }\] \[m = undefined. \]

OpenStudy (anonymous):

The slope is undefined, but the equation of the line can still be written.

OpenStudy (anonymous):

how would it look like if its undefined?

OpenStudy (anonymous):

@Brittni0605 I never said you could not do that, the user who asked this was offline when I was helping him. But as the slope of the y-axis is undefined the equation must be of the form x=k

OpenStudy (anonymous):

so x=k or just simply undefined?

OpenStudy (anonymous):

y = mx+c; if slope undefined this imply m tends to infinity; line is parallel to y axis given that the line passes through

OpenStudy (anonymous):

soo??